OpenStudy (anonymous):

I NEED MAJOR HELP IM TRYING TO SOLVE EQUATIONS WITH EXPONENTS (X+1)^3/2=64

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

x^3 + 1 /2 = 64 Find the greatest common denominator, so 2 x 64 = 128 x^3 + 1 = 128 ( cancel out denominators) x^3 = 127

OpenStudy (anonymous):

I may be wrong however

OpenStudy (anonymous):

yeah, you are wrong.. :)

OpenStudy (anonymous):

lol thought so

OpenStudy (anonymous):

multiply both sides by 2. (x+1)^3 = 128. to get rid of the exponents, take the cube root of both sides. x+1 = cbrt (128). x = cbrt(128) - 1

OpenStudy (anonymous):

that's assuming the 2 is not part of the exponent

OpenStudy (anonymous):

Ahh nicely done man I'm only 15 so forgot to cube root :P

OpenStudy (anonymous):

no problem i'm older than that be careful about giving out help if you're not sure though :)

OpenStudy (anonymous):

SO WHATS THE ANSWER A)15B)511C)17D)21

OpenStudy (amistre64):

we should verify what the equation actually is because the way it is is confusing.

OpenStudy (amistre64):

(X+1)^(3/2)=64 ? or [(X+1)^3]/2=64

OpenStudy (anonymous):

(X+1)^2/3=64

OpenStudy (amistre64):

thats what you had typed before, and it is ambiguous. Pick one of the ones I offered please.

OpenStudy (anonymous):

THE FIRST

OpenStudy (amistre64):

ok... (X+1)^(3/2)=64 ln((x+1)^(3/2)) = ln(64) (3/2) ln(x+1) = ln(64) log2(x+1) = (2/3) log2(64) log2(x+1) = (2/3) (6) log2(x+1) = 4 x+1 = 2^4 x+1 = 16 x = 15 if I did it all right :)

OpenStudy (anonymous):

THANKS SO MUCH